# A Comprehensive Overview of Large Language Models

### v9

@@ 5.2 Evaluation Datasets and Tasks
Natural Language Understanding: This task measures the language understanding capacity of LMs. It encompasses multiple tasks, including sentiment analysis, text classification, natural language inference (NLI), question answering (QA), commonsense reasoning (CR), mathematical reasoning (MR), reading comprehension (RC), etc.
Natural Language Generation: This task assesses the language generation capabilities of LLMs by understanding the provided input context. It includes tasks such as summarization, sentence completion, machine translation (MT), dialogue generation, etc.
@@
Anthropic-$H$-RLHF [296] Alignment 142k - - Manual
Anthropic-$H$-RLHF-2 [168] Alignment 39k - - Manual
@@
ANLI [383], MNLI-m [384], MNLI-$m$ [384],QNLI [352], WNLI [347], OCNLI [301], CMNLI [301], ANLI R1 [383], ANLI R2 [383], ANLI R3 [383], HANS [385], OCNLI-FC [302], LogiQA [386], StrategyQA [339]
@@ 6 Applications
General Purpose: LLMs are being widely considered as general-purpose tools for a wide variety of tasks [421]. This is due to their inherent ability to understand, generate, and manipulate human-like text in a contextually relevant manner. This allows them to perform tasks ranging from simple language translation and question-answering to more complex tasks like summarization, text generation, and even programming help [422]. The utility of LLMs is further enhanced by their ability to adapt to the specific style and tone of the text they are processing, making the outputs more user-friendly and context-aware. In everyday applications, LLMs can be used as personal assistants, helping users draft emails or schedule appointments [423]; they can also be deployed in customer service to handle common questions; or applied to generate content for digital platforms like websites, by creating human-like text based on given prompts [424]. Moreover, LLMs play a crucial role in data analysis, where they can filter large volumes of text data, summarize key points, and find patterns that would take humans much longer to identify [425]. Despite their wide-ranging applications, it is essential to remember that LLMs, similar to any AI system, are only as good as the data they have been trained on.
@@ 7 Challenges and Future Directions
Computational Cost: Training LLMs requires extensive computational resources, which increases production costs and raises environmental concerns due to substantial energy consumption during large-scale training. Improved performance occurs as computational resources increase, but the rate of improvement gradually decreases when both the model and dataset size remain fixed, following the power law of diminishing returns [467].
@@
Hallucinations: LLMs exhibit “hallucinations", where they generate responses that, while sounding plausible, are incorrect or do not align with the provided information [473]. The hallucination can be categorized into three categories.
@@
Limited Knowledge: Information acquired during pretraining is limited and may become obsolete after some time. Re-training the model using updated data is costly. To generate factually accurate responses people use a retrieval augmentation pipeline [188]. However, pre-trained models are not trained with retrieval augmentation generation (RAG) [6, 21], hence, adapting the training pipeline is necessary [183, 25].
@@
Catastrophic Forgetting: LLMs are often pre-trained on large datasets and then fine-tuned on domain-specific data, reducing training resources but facing issues like domain adaptation and catastrophic forgetting, which hinders the retention of original knowledge when learning new tasks.
@@
Interpretability and Explainability: The "black-box" nature of LLMs poses challenges in understanding their decision-making, which is crucial for broader acceptance and trust, especially in sensitive domains. Despite their advanced capabilities, the lack of insight into their operation limits their effectiveness and trustworthiness [479, 480]. Efforts are being made to make LLMs more explainable to promote user trust and to ensure responsible AI usage. Understanding the logic behind LLMs’ responses is essential for fostering trust and ensuring they align with human values and legal standards.
@@
Long-Term Dependencies: Large Language Models (LLMs) have shown considerable progress in understanding and generating text, yet they often struggle with preserving context and handling long-term dependencies, particularly in complex, multi-turn conversations or long documents. This limitation can lead to incoherent or irrelevant responses.
@@ 8 Conclusion
This article has reviewed the developments on LLMs comprehensively. It contributes to summarizing significant findings of LLMs in the existing literature and provides a detailed analysis of the design aspects, including architectures, datasets, and training pipelines. We identified crucial architectural components and training strategies employed by different LLMs. These aspects are presented as summaries and discussions throughout the article. Moreover, we have discussed the performance differences of LLMs in zero-shot and few-shot settings, explored the impact of fine-tuning, and compared supervised and generalized models and encoder vs. decoder vs. encoder-decoder architectures. A comprehensive review of multi-modal LLMs, retrieval augmented LLMs, LLMs-powered agents, efficient LLMs, datasets, evaluation, applications, and challenges is also provided. This article is anticipated to serve as a valuable resource for researchers, offering insights into the recent advancements in LLMs and providing fundamental concepts and details to develop better LLMs.

### v10

@@ 5.2 Evaluation Datasets and Tasks
  Natural Language Understanding: It measures the language understanding capacity of LMs. It encompasses multiple tasks, including sentiment analysis, text classification, natural language inference (NLI), question answering (QA), commonsense reasoning (CR), mathematical reasoning (MR), reading comprehension (RC), etc.
  Natural Language Generation: It assesses the language generation capabilities of LLMs by understanding the provided input context. It includes tasks such as summarization, sentence completion, machine translation (MT), dialogue generation, etc.
@@
Anthropic-HH-RLHF [306] Alignment 142k - - Manual
Anthropic-HH-RLHF-2 [178] Alignment 39k - - Manual
@@
ANLI [393], MNLI-m [394], MNLI-mm [394],QNLI [362], WNLI [357], OCNLI [311], CMNLI [311], ANLI R1 [393], ANLI R2 [393], ANLI R3 [393], HANS [395], OCNLI-FC [312], LogiQA [396], StrategyQA [349]
@@ 6 Applications
General Purpose: LLMs are being widely considered as general-purpose tools for a wide variety of tasks [431]. This is due to their inherent ability to understand, generate, and manipulate human-like text in a contextually relevant manner. This allows them to perform tasks ranging from simple language translation and question-answering to more complex tasks like summarization, text generation, and even programming help [432]. The utility of LLMs is further enhanced by their ability to adapt to the specific style and tone of the text they are processing, making the outputs more user-friendly and context-aware. In everyday applications, LLMs can be used as personal assistants, helping users draft emails or schedule appointments [433]; they can also be deployed in customer service to handle common questions or applied to generate content for digital platforms like websites by creating human-like text based on given prompts [434]. Moreover, LLMs play a crucial role in data analysis, where they can filter large volumes of text data, summarize key points, and find patterns that would take humans much longer to identify [435]. Despite their wide-ranging applications, it is essential to remember that LLMs, similar to any AI system, are only as good as the data they have been trained on.
@@ 7 Challenges and Future Directions
Computational Cost: Training LLMs require extensive computational resources, which increases production costs and raises environmental concerns due to substantial energy consumption during large-scale training. Improved performance occurs as computational resources increase, but the rate of improvement gradually decreases when both the model and dataset size remain fixed, following the power law of diminishing returns [477].
@@
Hallucinations: LLMs exhibit “hallucinations", where they generate responses that, while sounding plausible, are incorrect or do not align with the provided information [483]. Hallucinations can be categorized into three categories.
@@
Limited Knowledge: Information acquired during pretraining is limited and may become obsolete after some time. Re-training the model using updated data is costly. To generate factually accurate responses, people use a retrieval augmentation pipeline [198]. However, pre-trained models are not trained with retrieval augmentation generation (RAG) [6, 21]; hence, adapting the training pipeline is necessary [193, 25].
@@
Security and Privacy: LLMs are prone to leaking personal information and generating false, unethical, misaligned responses. Researchers have explored various security attacks, i.e., backdoor attacks, jailbreaking, prompt injection, and data poisoning, that lead to breaking LLMs security. Therefore, developing better defense mechanisms is essential to ensure LLMs are safe, reliable, and trustworthy for complex AI applications [486].
@@
Catastrophic Forgetting: LLMs are often pre-trained on large datasets and then fine-tuned on domain-specific data, reducing training resources. However, they face issues like domain adaptation and catastrophic forgetting, which hinder the retention of original knowledge when learning new tasks.
@@
Interpretability and Explainability: The “black-box” nature of LLMs poses challenges in understanding their decision-making, which is crucial for broader acceptance and trust, especially in sensitive domains. Despite their advanced capabilities, the lack of insight into their operation limits their effectiveness and trustworthiness [490, 491]. Efforts are being made to make LLMs more explainable to promote user trust and to ensure responsible AI usage. Understanding the logic behind LLMs’ responses is essential for fostering trust and ensuring they align with human values and legal standards.
@@
Long-Term Dependencies: Large Language Models have shown considerable progress in understanding and generating text, yet they often struggle with preserving context and handling long-term dependencies, particularly in complex, multi-turn conversations or long documents. This limitation can lead to incoherent or irrelevant responses.
@@ 8 Conclusion
This article has comprehensively reviewed the developments in LLMs. It contributes to summarizing significant findings of LLMs in the existing literature and provides a detailed analysis of the design aspects, including architectures, datasets, and training pipelines. We identified crucial architectural components and training strategies employed by different LLMs. These aspects are presented as summaries and discussions throughout the article. Moreover, we have discussed the performance differences of LLMs in zero-shot and few-shot settings, explored the impact of fine-tuning, and compared supervised and generalized models and encoder vs. decoder vs. encoder-decoder architectures. A comprehensive review of multi-modal LLMs, retrieval augmented LLMs, LLMs-powered agents, efficient LLMs, datasets, evaluation, applications, and challenges is also provided. This article is anticipated to serve as a valuable resource for researchers, offering insights into the recent advancements in LLMs and providing fundamental concepts and details to develop better LLMs.
Acknowledgement: The author/s would like to acknowledge the support received from Saudi Data and AI Authority (SDAIA) and King Fahd University of Petroleum and Minerals (KFUPM) under SDAIA-KFUPM Joint Research Center for Artificial Intelligence Grant No. JRC-AI-RFP-11.